# Anytime-Constrained Multi-Agent Reinforcement Learning

### v1

@@ Abstract
We introduce anytime constraints to the multi-agent setting with the corresponding solution concept being anytime-constrained equilibrium (ACE). Then, we present a comprehensive theory of anytime-constrained Markov games, which includes (1) a computational characterization of feasible policies, (2) a fixed-parameter tractable algorithm for computing ACE, and (3) a polynomial-time algorithm for approximately computing feasible ACE. Since computing a feasible policy is NP-hard even for two-player zero-sum games, our approximation guarantees are the best possible under worst-case analysis. We also develop the first theory of efficient computation for action-constrained Markov games, which may be of independent interest.
@@ 1 Introduction
Although multi-agent reinforcement learning (MARL) has made many breakthroughs in game-playing, the literature has long since advocated the importance of constraints in real-world applications [22]. Despite their importance, the literature on constrained MARL is far behind the state-of-the-art in the single-agent setting. Most recently, almost-sure [8] and anytime [30, 29] constraints have emerged in the single-agent setting to capture real-world scenarios, such as medical applications [13, 31, 25], disaster relief scenarios [18, 43, 38], and resource management [27, 26, 33, 4]. Anytime constraints are especially natural for autonomous vehicles [30, 36, 43], which are a primary application of MARL [12, 15, 42]. For example, a GPS client must balance safety with agents’ selfish desires lest they risk accidents that would dissuade agents from using the service. Despite these applications, anytime constraints have yet to be studied in the multi-agent setting, which we remedy in this work.
@@ 1.1 Related Work
Most works implement safety using some constraints [19], and several multi-agent works exist down this line [36, 21, 17]. The single agent setting has mainly focused on no-violation learning for cMDPs [11, 6, 22] and solving CCMDPs [40, 22], which capture the probability of entering unsafe states. Performing learning while avoiding dangerous states has also been studied [35, 37, 44] under non-trivial assumptions.
@@ 2 Equilibria
Solutions to games traditionally take the form of equilibrium. In the MARL realm, the most popular notions include the Nash equilibrium (NE), correlated equilibrium (CE), and course-correlated equilibrium (CCE). Given constraints, the key difference is a focus on feasible policies. Infeasible policies lead to disastrous outcomes for an agent. Thus, not only should a constrained equilibrium be feasible, but agents should only consider deviating if doing so would be feasible.
@@
Our definition of ACE in Definition 1 technically corresponds to anytime-constrained course-correlated equilibria (ACCCE), which we simplify to ACE for exposition purposes. Our results apply equally well to anytime-constrained correlated equilibria (ACCE). We delay the definition and discussion of ACCE to the appendix. In the main text, we signal when results specialize to each equilibria type by writing ACE (NE/CE/CCE).
@@
It is often useful to consider refinements of equilibrium notions that are more structured and robust. The classical refinement for sequential games is the subgame-perfect equilibrium (SPE). An SPE policy is required to behave optimally under any history, also called subgames, even if those subgames are not realizable. That way, players could still recover if any deviated from the policy’s suggestion. In the constrained setting, we only consider feasible subgames, i.e., subgames that are realizable by some feasible policy. This means the players could still adapt and finish the game whenever a player takes an unsupported but feasible action.
@@
denote the set of feasible policies for the subgame $\tau_{h}$. To capture our earlier intuition, we require an anytime-constrained SPE to be a policy that is feasible for any feasible subgame, and that beats any feasible deviation for that subgame.
@@
Although Proposition 1 implies equilibria exist under very minimal assumptions, they are generally hard to compute. As shown in  [30], feasible policies are generally not Markovian nor product policies. Moreover, just determining whether there exists a feasible policy is NP-hard even for the simplest of acMGs.
@@ 3 Feasibility
Before we can fully understand ACE, we must first understand feasible policies. This section derives characterizations of feasibly realizable histories under anytime constraints. This leads us to design an algorithm that determines if a feasible policy exists, while also producing the set of all feasibly realizable cumulative costs and actions. These sets will be critical to our later equilibria computation in Section 4.
@@
Our graph is generated by iteratively taking feasible actions. If no sequence of feasible actions ever reaches time $H+1$, then naturally no feasible policy exists. However, we must also ensure all branches generated from an action reach time $H+1$ to satisfy the anytime constraints. We deal with this difficulty by making each action an AND node. On the other hand, for a (time, state, cost)-triple to be feasible, there need only be a single action that ensures time $H+1$ is reached. Thus, we make each triple an OR node. We will later show that a TRUE subgraph corresponds to our desired solution.
@@
Overall, we see the subtree of TRUE nodes of $\mathcal{T}$ is exactly the union of all feasible policy realizability graphs. Computing the TRUE nodes for an AND/OR tree can be done in linear time using standard tree recursion [28]. Suppose that AOSolve is any such AND/OR tree solver. Then, we can compute the $\mathcal{FS}$ and $\mathcal{FA}$ sets by implicitly pruning the FALSE nodes from $\mathcal{T}$. The full procedure is described in Algorithm 1.
@@ 4 Reduction
As hinted in the previous section, we can convert the anytime constraint on full histories into a per-time constraint on the available actions. Specifically, if the agents track their cumulative costs, they can identify actions that satisfy the constraint long term. These actions exactly correspond to those in $\mathcal{FA}_{h}(s,\bar{c})$.
@@
For typical MGs, the standard solution concept is the Markov-perfect equilibrium. For action-constrained MGs, we use the same solution concept, but add the condition that the policy must only support actions in the constrained action sets.
@@ 5 Computation
In the last section, we showed how to reduce our anytime-constrained game problem to an action-constrained game problem. However, efficient algorithms for action-constrained games are currently unknown. In this section, we remedy this knowledge gap by designing efficient algorithms for computing MPE of an action-constrained MG. Moreover, we show that feeding our action-constrained method into our reduction yields a polynomial time algorithm for computing ACE so long as the cost precision is logarithmic.
@@
The key bottleneck to this backward induction approach is solving the action-constrained matrix games. Formally, let $(\mathcal{A},X,u)$ denote an action-constrained matrix game, where (i) $\mathcal{A}$ is the joint action space, (ii) $X\subseteq\mathcal{A}$ is the set of feasible actions, and (iii) $u$ is the utility function. We tackle this problem by devising a variation of the standard CE/CCE LP. Importantly, we modify the constraint $\sum_{a\in\mathcal{A}}\sigma(a)=1$, which ensures the total probability mass of all joint actions equals one, into the constraint $\sum_{a\in X}\sigma(a)=1$, which ensures the support of the joint strategy is contained in the valid joint action space. We also define the utility of any infeasible action to be $\infty$- ∞ so that infeasible deviations will be appropriately ignored by the LP. The full definition of the LP, which has no objective function, is,
@@
Given our efficient action-constrained game solver, we can now finish analyzing the running time of our reduction. Since Theorem 2 implies that Algorithm 3 runs in time ${\rm poly}(|\overline{G}|)$, Theorem 1 implies that Algorithm 2 with MGSolve = Algorithm 3 runs in time ${\rm poly}(|G|,D_{G})$. The main issue is that $D_{G}$ can be exponentially large in the worst case. However, we can show that $D_{G}\leq{\rm poly}(|G|)2^{O(dn)}$, where $d$ denotes the cost precision, which is the number of significant bits needed to represent any supported cost. By definition, our method is FPT [16] in the cost precision.
@@ 6 Approximation
In the previous section, we showed our method runs in polynomial time whenever the cost precision is small. However, in cases where the cost precision is large or, even worse infinite, the computation may require exponential time due to the NP-hard nature of finding a feasible solution, Proposition 2. To combat this issue, we slightly relax the feasibility condition. This allows us to compute equilibrium policies that only violate the budget by a given $\epsilon>0$ at the cost of an additional ${\rm poly}(1/\epsilon)$ factor in running time.
In this section, we allow any infinite support cost distributions that are bounded above. We also require that distribution is a product distribution to enable comparisons between supported costs. Technically, we also need the CDF of the distribution to be efficiently computable for use in computation.
@@
Moreover, if $Hc^{max}_{i}\leq B$, observe that every policy is feasible for player $i$, which just leads to a standard unconstrained problem for that player. A similar phenomenon happens if $c^{max}_{i}\leq 0$. Thus, we assume WLOG that $Hc^{max}>B$ and $c^{max}>0$. We can then define our relaxed feasibility notions as follows.
@@
The key idea of our approximations is to have players round down any cost vector they receive to the nearest multiple of some $\ell>0$. Doing so allows the players to track far fewer cumulative costs than originally. In addition, we can effectively truncate the lower regime of the distribution using the fact that if player $i$ ever receives an immediate cost smaller than $B_{i}-Hc^{max}_{i}$, then it may take any action whatsoever going forward without violating its budget. Thus, the player can treat any smaller cost as if it were $B_{i}-Hc^{max}_{i}$. This process of rounding costs then induces a new cMG with finite support cost distributions.
@@
We then define the approximate acMG $\hat{G}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{S},\mathcal{A},P,R,\hat{C},H)$ to be our original game $G$ but with a different cost distribution.
Since outside of extreme cases, we always round costs down, the players always maintain an underestimate of their true cumulative cost. Consequently, the players may be incurring more costs than expected. However, we can show that the true cost accumulated is not much larger than the surrogate cost.
@@
On the flip side, the fact that players are willing to spend more than before means they have more actions available to them at each stage. Consequently, they can always find strategies that achieve the same value or even higher than any truly feasible policy for the game. It is then easy to see that solutions to $\hat{G}$ satisfy condition (2) in Definition 1, so form approximate ACE for $G$.
@@ 7 Conclusion
In this work, we introduced anytime-constraints for Markov games and studied the corresponding solution concept of anytime-constrained equilibria. Although finding a feasible policy is NP-hard for simple games, we showed efficient computation is possible so long as the cost precision is constant. The main ingredients to our approach were a graph algorithm to derive all feasibly realizable histories and an efficient algorithm for solving action-constrained MGs. Lastly, we presented approximation algorithms for computing approximately feasible anytime-constrained equilibria running in polynomial time so long as the cost distribution’s supremum is no larger than a polynomial factor of the budget. Given the hardness results, our approximation guarantees are best possible under worst-case analysis.

### v2

# Anytime-Constrained Equilibria in Polynomial Time

@@ Abstract
We extend anytime constraints to the Markov game setting and the corresponding solution concept of an anytime-constrained equilibrium (ACE). Then, we present a comprehensive theory of anytime-constrained equilibria that includes (1) a computational characterization of feasible policies, (2) a fixed-parameter tractable algorithm for computing ACE, and (3) a polynomial-time algorithm for approximately computing ACE. Since computing a feasible policy is NP-hard even for two-player zero-sum games, our approximation guarantees are optimal so long as $P\neq NP$. We also develop the first theory of efficient computation for action-constrained Markov games, which may be of independent interest.
@@ 1 Introduction
Although multi-agent reinforcement learning (MARL) has made many breakthroughs in game-playing, the literature has long since advocated the importance of constraints in real-world applications [22]. Despite their importance, the literature on constrained MARL is far behind the state-of-the-art in the single-agent setting. Most recently, almost-sure [8] and anytime [30, 29] constraints have emerged in the single-agent setting to capture real-world scenarios, such as medical applications [13, 31, 25], disaster relief scenarios [18, 43, 38], and resource management [27, 26, 33, 4]. However, many of these motivating applications are actually multi-agent problems. Most obviously, anytime-compliant autonomous vehicles [30, 36, 43] must interact with other vehicles, which is a key aspect of MARL [12, 15, 42]. Despite their relevance, anytime constraints have yet to be studied in the multi-agent setting, which we remedy in this work.
@@ 1.1 Related Work
Most works implement safety using some constraints [19], and several multi-agent works exist down this line [36, 21, 17]. The single-agent setting has mainly focused on no-violation learning for cMDPs [11, 6, 22] and solving CCMDPs [40, 22], which capture the probability of entering unsafe states. Performing learning while avoiding dangerous states has also been studied [35, 37, 44] under non-trivial assumptions.
@@ 2 Equilibria
Solutions to games traditionally take the form of equilibrium. In the MARL realm, the most popular notions include the Nash equilibrium (NE), correlated equilibrium (CE), and course-correlated equilibrium (CCE). Given the constraints, the key difference is a focus on feasible policies. Infeasible policies lead to disastrous outcomes for an agent. Thus, not only should a constrained equilibrium be feasible, but agents should only consider deviating if doing so would be feasible.
@@
Our definition of ACE in Definition 1 technically corresponds to anytime-constrained course-correlated equilibria (ACCCE), which we simplify to ACE for exposition purposes. Our results apply equally well to anytime-constrained correlated equilibria (ACCE). We delay the definition and discussion of ACCE to the appendix. In the main text, we signal when results specialize to each equilibrium type by writing ACE (NE/CE/CCE).
@@
It is often useful to consider refinements of equilibrium notions that are more structured and robust. The classical refinement for sequential games is the subgame-perfect equilibrium (SPE). An SPE policy is defined to behave optimally under any history, also called subgames, even if those subgames are not realizable. That way, players could still recover if they deviated from the policy’s suggestion. In the constrained setting, we only consider feasible subgames, i.e., subgames that are realizable by some feasible policy. This means the players could still adapt and finish the game whenever a player takes an unsupported but feasible action.
@@
denote the set of feasible policies for the subgame $\tau_{h}$. To capture our earlier intuition, we require an anytime-constrained SPE to be a feasible policy for any feasible subgame and that beats any feasible deviation for that subgame.
@@
Although Proposition 1 implies equilibria exist under minimal assumptions, they are generally hard to compute. As shown in  [30], feasible policies are generally not Markovian nor product policies. Moreover, just determining whether a feasible policy exists is NP-hard, even for the simplest of acMGs.
@@ 3 Feasibility
Before we can fully understand ACE, we must first understand feasible policies. This section derives characterizations of feasibly realizable histories under anytime constraints. This leads us to design an algorithm that determines if a feasible policy exists while also producing the set of all feasibly realizable cumulative costs and actions. These sets will be critical to our later equilibria computation in Section 4.
@@
Our graph is generated by iteratively taking feasible actions. If no sequence of feasible actions ever reaches time $H+1$, then naturally no feasible policy exists. However, we must also ensure that all branches generated from an action reach time $H+1$ to satisfy the anytime constraints. We deal with this difficulty by making each action an AND node. On the other hand, for a (time, state, cost)-triple to be feasible, there needs only be a single action that ensures time $H+1$ is reached. Thus, we make each triple an OR node. We will later show that a TRUE subgraph corresponds to our desired solution.
@@
Overall, the subtree of TRUE nodes of $\mathcal{T}$ is exactly the union of all feasible policy realizability graphs. Computing the TRUE nodes for an AND/OR tree can be done in linear time using standard tree recursion [28]. Suppose that AOSolve is any such AND/OR tree solver. Then, we can compute the $\mathcal{FS}$ and $\mathcal{FA}$ sets by implicitly pruning the FALSE nodes from $\mathcal{T}$. The full procedure is described in Algorithm 1.
@@ 4 Reduction
As hinted in the previous section, we can convert the anytime constraint on full histories into a per-time constraint on the available actions. Specifically, if the agents track their cumulative costs, they can identify actions that satisfy the constraint in the long term. These actions exactly correspond to those in $\mathcal{FA}_{h}(s,\bar{c})$.
@@
For typical MGs, the standard solution concept is the Markov-perfect equilibrium. For action-constrained MGs, we use the same solution concept but add the condition that the policy must only support actions in the constrained action sets.
@@ 5 Computation
In the last section, we showed how to reduce our anytime-constrained game problem to an action-constrained one. However, efficient algorithms for action-constrained games are currently unknown. In this section, we remedy this knowledge gap by designing efficient algorithms for computing an MPE of an action-constrained MG. Moreover, we show that feeding our action-constrained method into our reduction yields a polynomial time algorithm for computing ACE so long as the cost precision is logarithmic.
@@
The key bottleneck to this backward induction approach is solving the action-constrained matrix games. Formally, let $(\mathcal{A},X,u)$ denote an action-constrained matrix game, where (i) $\mathcal{A}$ is the joint action space, (ii) $X\subseteq\mathcal{A}$ is the set of feasible actions, and (iii) $u$ is the utility function. We tackle this problem by devising a variation of the standard CE/CCE LP. Importantly, we modify the constraint $\sum_{a\in\mathcal{A}}\sigma(a)=1$, which ensures the total probability mass of all joint actions equals one, into the constraint $\sum_{a\in X}\sigma(a)=1$, which ensures the support of the joint strategy is contained in the valid joint action space. We also define the utility of any infeasible action to be $\infty$- ∞ so that the LP will appropriately ignore infeasible deviations. The full definition of the LP, which has no objective function, is,
@@
Given our efficient action-constrained game solver, we can now finish analyzing the run time of our reduction. Since Theorem 2 implies that Algorithm 3 runs in time ${\rm poly}(|\overline{G}|)$, Theorem 1 implies that Algorithm 2 with MGSolve = Algorithm 3 runs in time ${\rm poly}(|G|,D_{G})$. The main issue is that $D_{G}$ can be exponentially large in the worst case. However, we can show that $D_{G}\leq{\rm poly}(|G|)2^{O(dn)}$, where $d$ denotes the cost precision, which is the number of significant bits needed to represent any supported cost. By definition, our method is FPT [16] in the cost precision.
@@ 6 Approximation
In the previous section, we showed our method runs in polynomial time whenever the cost precision is small. However, in cases where the cost precision is large or, even worse, infinite, the computation may require exponential time due to the NP-hard nature of finding a feasible solution, Proposition 2. To combat this issue, we slightly relax the feasibility condition. This allows us to compute equilibrium policies that only violate the budget by a given $\epsilon>0$ at the cost of an additional ${\rm poly}(1/\epsilon)$ factor in running time.
In this section, we allow any infinite support cost distributions that are bounded above. We also require that distribution be a product distribution to enable comparisons between supported costs. Technically, we also need the CDF of the distribution to be efficiently computable for use in computation.
@@
Moreover, if $Hc^{max}_{i}\leq B$, observe that every policy is feasible for player $i$, which leads to a standard unconstrained problem for that player. A similar phenomenon happens if $c^{max}_{i}\leq 0$. Thus, we assume WLOG that $Hc^{max}>B$ and $c^{max}>0$. We can then define our relaxed feasibility notions as follows.
@@
The key idea of our approximations is to have players round down any cost vector they receive to the nearest multiple of some $\ell>0$. Doing so allows the players to track far fewer cumulative costs than originally. In addition, we can effectively truncate the lower regime of the distribution using the fact that if player $i$ ever receives an immediate cost smaller than $B_{i}-Hc^{max}_{i}$, then it may take any action going forward without violating its budget. Thus, the player can treat any smaller cost as if it were $B_{i}-Hc^{max}_{i}$. This rounding process then induces a new cMG with finite support cost distributions.
@@
We then define the approximate acMG $\hat{G}\stackrel{{\scriptstyle\text{def}}}{{=}}(\mathcal{S},\mathcal{A},P,R,\hat{C},H)$ as our original game $G$ but with a different cost distribution.
Since we always round costs down except in extreme cases, the players always underestimate their true cumulative cost. Consequently, the players may be incurring more costs than expected. However, we can show that the true cost accumulated is not much larger than the surrogate cost.
@@
On the flip side, the fact that players are willing to spend more than before means they have more actions available at each stage. Consequently, they can always find strategies that achieve the same value or even higher than any truly feasible policy for the game. It is then easy to see that solutions to $\hat{G}$ satisfy condition (2) in Definition 1, so form an approximate ACE for $G$.
@@ 7 Conclusion
In this work, we introduced anytime-constraints for Markov games and studied the corresponding solution concept of anytime-constrained equilibria. Although finding a feasible policy is NP-hard for simple games, we showed efficient computation is possible so long as the cost precision is constant. The main ingredients of our approach were a graph algorithm to derive all feasibly realizable histories and an efficient algorithm for solving action-constrained MGs. Lastly, we presented approximation algorithms for computing approximately feasible anytime-constrained equilibria running in polynomial time so long as the cost distribution’s supremum is no larger than a polynomial factor of the budget. Given the hardness results, our approximation guarantees are the best possible under worst-case analysis.